OpenStudy (anonymous):

1. The relationship 1/sin(θ) is equivalent to which of the following trigonometric functions? A. cot(θ) B. csc(θ) C. arcsin(θ) D. sec(θ)

hartnn (hartnn):

sin x = 1/csc x so what will be 1/sin x ?

OpenStudy (anonymous):

idk :/

hartnn (hartnn):

if a = 1/b then b = 1/a !! so if sinx = 1/csc x then 1/ sin x = csc x :)

hartnn (hartnn):

here you just have \(\theta \) instead of x ..